a 15% solution of alcohol is to be mixed with a 40% solution to get 50 liters of a final mixture that is 30% alcohol. how much of each of the oiginal solutions should be used?
Answer by mananth(16946)

a 15% solution of alcohol is to be mixed with a 40% solution to get 50 liters of a final mixture that is 30% alcohol. how much of each of the oiginal solutions should be used?
...
15% solution of alcohol -----x liters
40% solution ----------------50-x liters
to get 50 liters mixture that is 30% alcohol.
...
0.15x+0.4(50-x)=0.3*50
0.15x+20-0.4x=15
0.25x=-20+15
0.25x=-5
x=-5/0.25
x=20 liters
...
m.ananth@hotmail.ca